# Higher Lie characters and root enumeration in classical Weyl groups

### v1

@@ 4.1. Definitions
    It is easy to see that $\psi^{x}$ depends only on the conjugacy class $C$ (equivalently, the cycle type ${\underline{\lambda}}$) of $x$, and can therefore be denoted $\psi^{C}$ or $\psi^{\underline{\lambda}}$.
@@ 4.2. Main result
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\frac{\psi^{\underline{\lambda}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$ $=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}\frac{K_{\varepsilon,\theta}(e)(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@ 4.3. Values of induced characters
For an element $x\in B_{n}$ denote the conjugacy class and the centralizer of $x$ by $C_{x}$ and $Z_{x}$, respectively. Recall the definitions of $\omega^{x}$ and $\psi^{x}$ from Subsection 4.1.
@@
$\psi^{x}(g)=\frac{|C_{x}|}{|C_{g}|}\sum_{z\in C_{g}\cap Z_{x}}\omega^{x}(z)\qquad(\forall g\in B_{n}).$
@@
$\psi^{x}(g)=\omega^{x}\uparrow_{Z_{x}}^{B_{n}}(g)=\frac{|Z_{g}|}{|Z_{x}|}\sum_{z\in C_{g}\cap Z_{x}}\omega^{x}(z)\qquad(\forall g\in B_{n}).$
@@ 4.4. The structure of a single cycle
      with $p_{1,t+k}$ interpreted as $-p_{1,t+k-i}$ if $t+k>i$ (thus as as $p_{1,t+k-2i}$ if $t+k>2i$).
@@ 4.5. Summation on a single cycle
Extending the previous result, we shall now sum $\omega^{x}$ on the whole set $C_{g}\cap Z_{x}$, with increasing generality of the cycle types of $g$ and $x$. This will lead to a proof of Theorem 4.3, providing a generating function for the values of higher Lie characters $\psi^{\underline{\lambda}}=\psi^{x}$.
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon,j,\theta}(m)}\omega^{x}(z)\frac{s^{m}}{m!}=\exp\left(\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(2is)^{j/e}}{2ij/e}\right),$
@@
$=m!\sum_{(n_{1},\ldots,n_{q})\in N_{i,j}(m)}\prod_{k=1}^{q}\frac{1}{n_{k}!}\left(\frac{K_{\varepsilon,\theta}(e_{k})(2i)^{\ell_{k}}}{2i\ell_{k}}\right)^{n_{k}}.$
@@
$\prod_{k=1}^{q}\sum_{n_{k}=0}^{\infty}\frac{1}{n_{k}!}\left(\frac{K_{\varepsilon,\theta}(e_{k})(2is)^{\ell_{k}}}{2i\ell_{k}}\right)^{n_{k}}$ $=\prod_{k=1}^{q}\exp\left(\frac{K_{\varepsilon,\theta}(e_{k})(2is)^{\ell_{k}}}{2i\ell_{k}}\right)$
$=\prod_{e\in E(i,j)}\exp\left(\frac{K_{\varepsilon,\theta}(e)(2is)^{j/e}}{2ij/e}\right).$
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon,j,\theta}(m)}\omega^{x}(z)\frac{s^{m}}{m!}$ $=\prod_{e\in E(i,j)}\exp\left(\frac{K_{\varepsilon,\theta}(e)(2is)^{j/e}}{2ij/e}\right)$
$=\exp\left(\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(2is)^{j/e}}{2ij/e}\right).\qed$
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon}(m)}\omega^{x}(z)\frac{s^{m}}{m!}\prod_{j,\theta}t_{j,\theta}^{jm_{j,\theta}(z)}=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right).$
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon}(m)}\omega^{x}(z)\frac{s^{m}}{m!}\prod_{j,\theta}t_{j,\theta}^{jm_{j,\theta}(z)}$ $=\prod_{j,\theta}\exp\left(\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right)$
$=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right),$
@@
Recalling the higher Lie characters $\psi^{\underline{\lambda}}$ from Definition 4.2(c), we can now prove the main result of this section, Theorem 4.3. It extends Lemma 4.12 and computes the sum in Lemma 4.4 on the whole set $C_{g}\cap Z_{x}$, for arbitrary $g$ and $x$.
@@
$\sum_{a_{i,\varepsilon}\geq 0}\sum_{z_{i,\varepsilon}\in G_{i,\varepsilon}\wr S_{a_{i,\varepsilon}}}\omega^{x_{i,\varepsilon}}(z_{i,\varepsilon})\frac{s^{a_{i,\varepsilon}}}{a_{i,\varepsilon}!}\prod_{j,\theta}t_{j,\theta}^{jm_{j,\theta}(z_{i,\varepsilon})}=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right).$
@@
$=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$=\prod_{i,\varepsilon}\Sigma_{i,\varepsilon}=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e\in E(i,j)}\frac{K_{\varepsilon,\theta}(e)(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$|C_{\underline{\lambda}}|\cdot\sum_{z\in Z_{x}\cap C_{\underline{\mu}}}\omega^{x}(z)=|C_{\underline{\mu}}|\cdot\psi^{x}({\underline{\mu}}).$
@@
LHS $=\sum_{n\geq 0}\frac{1}{|B_{n}|}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}|C_{\underline{\mu}}|\psi^{\underline{\lambda}}({\underline{\mu}}){\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$
$=\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\frac{\psi^{\underline{\lambda}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}.$
@@
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\frac{\psi^{\underline{\lambda}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$ $=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}\frac{K_{\varepsilon,\theta}(e)(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$\psi_{k}^{B_{n}}:=\sum_{{\underline{\lambda}}\vdash_{k}n}\psi^{\underline{\lambda}}.$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{\varphi_{k}^{B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j,\theta}^{jk/h}}{2jk/h}\cdot\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta)\right).$
@@
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\frac{\psi^{\underline{\lambda}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}\frac{K_{\varepsilon,\theta}(e)(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right),$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{\psi_{k}^{B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$ $=\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\sum_{{\underline{\lambda}}\vdash_{k}n}\frac{\psi^{\underline{\lambda}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$
$=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{ij/e}}{2ij/e}\right).$
@@
(1) $\sum_{\begin{subarray}{c}h\\ h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j,\theta}^{jk/h}}{2jk/h}\cdot\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta)=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{ij/e}}{2ij/e}.$
Letting $d:=i/e$ and then $h:=k/d$ in the RHS of (1) yields
RHS $=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|i\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=1\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{ij/e}}{2ij/e}$
$=\sum_{\begin{subarray}{c}d\\ d|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|(k/d)\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/(de)}=1\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{jd}}{2jd}$
$=\sum_{\begin{subarray}{c}h\\ h|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|h\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=1\end{subarray}}\frac{K_{\varepsilon,\theta}(e)t_{j,\theta}^{jk/h}}{2jk/h}.$
@@
(2) $\sum_{\begin{subarray}{c}e\\ e|\gcd(h,j)\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=1\end{subarray}}K_{\varepsilon,\theta}(e)=\begin{cases}\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta),&\text{if }\gcd(h,j)=1;\\ 0,&\text{otherwise.}\end{cases}$
@@
If $\gcd(h,j)\neq 1$ then the RHS here is zero, and if $\gcd(h,j)=1$ then the RHS is $\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta)$. This proves (2) and completes the proof of the theorem. ∎

### v2

@@ 4.1. Definitions
    It is easy to see that $\psi_{B_{n}}^{x}$ depends only on the conjugacy class $C$ (equivalently, the cycle type ${\underline{\lambda}}$) of $x$, and can therefore be denoted $\psi_{B_{n}}^{C}$ or $\psi_{B_{n}}^{\underline{\lambda}}$.
@@ 4.2. Main result
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,\frac{{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}$ $=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@ 4.3. Values of induced characters
For an element $x\in B_{n}$ denote the conjugacy class and the centralizer of $x$ by $C_{x}$ and $Z_{x}$, respectively. Recall the definitions of $\omega^{x}$ and $\psi_{B_{n}}^{x}$ from Subsection 4.1.
@@
$\psi_{B_{n}}^{x}(g)=\frac{|C_{x}|}{|C_{g}|}\sum_{z\in C_{g}\cap Z_{x}}\omega^{x}(z)\qquad(\forall g\in B_{n}).$
@@
$\psi_{B_{n}}^{x}(g)=\omega^{x}\uparrow_{Z_{x}}^{B_{n}}(g)=\frac{|Z_{g}|}{|Z_{x}|}\sum_{z\in C_{g}\cap Z_{x}}\omega^{x}(z)\qquad(\forall g\in B_{n}).$
@@ 4.4. The structure of a single cycle
      with $p_{1,t+k}$ interpreted as $-p_{1,t+k-i}$ if $t+k>i$ (thus as $p_{1,t+k-2i}$ if $t+k>2i$).
@@ 4.5. Summation on a single cycle
Extending the previous result, we shall now sum $\omega^{x}$ on the whole set $C_{g}\cap Z_{x}$, with increasing generality of the cycle types of $g$ and $x$. This will lead to a proof of Theorem 4.3, providing a generating function for the values of higher Lie characters $\psi_{B_{n}}^{\underline{\lambda}}=\psi_{B_{n}}^{x}$.
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon,j,\theta}(m)}\omega^{x}(z)\frac{s^{m}}{m!}=\exp\left(\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(2is)^{j/e}}{2ij/e}\right),$
@@
$=m!\sum_{(n_{1},\ldots,n_{q})\in N_{i,j}(m)}\prod_{k=1}^{q}\frac{1}{n_{k}!}\left(K_{\varepsilon,\theta}(e_{k})\,\frac{(2i)^{\ell_{k}}}{2i\ell_{k}}\right)^{n_{k}}.$
@@
$\prod_{k=1}^{q}\sum_{n_{k}=0}^{\infty}\frac{1}{n_{k}!}\left(K_{\varepsilon,\theta}(e_{k})\,\frac{(2is)^{\ell_{k}}}{2i\ell_{k}}\right)^{n_{k}}$ $=\prod_{k=1}^{q}\exp\left(K_{\varepsilon,\theta}(e_{k})\,\frac{(2is)^{\ell_{k}}}{2i\ell_{k}}\right)$
$=\prod_{e\in E(i,j)}\exp\left(K_{\varepsilon,\theta}(e)\,\frac{(2is)^{j/e}}{2ij/e}\right).$
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon,j,\theta}(m)}\omega^{x}(z)\,\frac{s^{m}}{m!}$ $=\prod_{e\in E(i,j)}\exp\left(K_{\varepsilon,\theta}(e)\,\frac{(2is)^{j/e}}{2ij/e}\right)$
$=\exp\left(\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(2is)^{j/e}}{2ij/e}\right).\qed$
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon}(m)}\omega^{x}(z)\,\frac{s^{m}}{m!}\prod_{j,\theta}t_{j,\theta}^{jm_{j,\theta}(z)}=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right).$
@@
$\sum_{m\geq 0}\sum_{z\in R_{i,\varepsilon}(m)}\omega^{x}(z)\,\frac{s^{m}}{m!}\prod_{j,\theta}t_{j,\theta}^{jm_{j,\theta}(z)}$ $=\prod_{j,\theta}\exp\left(\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right)$
$=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right),$
@@
Recalling the higher Lie characters $\psi_{B_{n}}^{\underline{\lambda}}$ from Definition 4.2(c), we can now prove the main result of this section, Theorem 4.3. It extends Lemma 4.12 and computes the sum in Lemma 4.4 on the whole set $C_{g}\cap Z_{x}$, for arbitrary $g$ and $x$.
@@
$\sum_{a_{i,\varepsilon}\geq 0}\sum_{z_{i,\varepsilon}\in G_{i,\varepsilon}\wr S_{a_{i,\varepsilon}}}\omega^{x_{i,\varepsilon}}(z_{i,\varepsilon})\,\frac{s^{a_{i,\varepsilon}}}{a_{i,\varepsilon}!}\prod_{j,\theta}t_{j,\theta}^{jm_{j,\theta}(z_{i,\varepsilon})}=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(2ist_{j,\theta}^{i})^{j/e}}{2ij/e}\right).$
@@
$=\exp\left(\sum_{j,\theta}\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$=\prod_{i,\varepsilon}\Sigma_{i,\varepsilon}=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e\in E(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$|C_{\underline{\lambda}}|\cdot\sum_{z\in Z_{x}\cap C_{\underline{\mu}}}\omega^{x}(z)=|C_{\underline{\mu}}|\cdot\psi_{B_{n}}^{x}({\underline{\mu}}).$
@@
LHS $=\sum_{n\geq 0}\frac{1}{|B_{n}|}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}|C_{\underline{\mu}}|\,\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}$
$=\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,\frac{{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}.$
@@
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,\frac{{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}$ $=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right).$
@@
$\psi_{k}^{B_{n}}:=\sum_{{\underline{\lambda}}\vdash_{k}n}\psi_{B_{n}}^{\underline{\lambda}}.$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\rho_{k}^{B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\rho_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}\right).$
@@
$\sum_{n\geq 0}\sum_{{\underline{\lambda}}\vdash n}\sum_{{\underline{\mu}}\vdash n}\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,\frac{{\underline{s}}^{{\underline{c}}({\underline{\lambda}})}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{i,\varepsilon}\sum_{j,\theta}\sum_{e|\gcd(i,j)}K_{\varepsilon,\theta}(e)\,\frac{(s_{i,\varepsilon}t_{j,\theta})^{ij/e}}{2ij/e}\right),$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\psi_{k}^{B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}$ $=\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\sum_{{\underline{\lambda}}\vdash_{k}n}\psi_{B_{n}}^{\underline{\lambda}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}$
$=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{ij/e}}{2ij/e}\right).$
@@
(1) $\sum_{\begin{subarray}{c}h\\ h|k\\ \gcd(h,j)=1\end{subarray}}\rho_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=1\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|\gcd(i,j)\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{ij/e}}{2ij/e}.$
Letting $d:=i/e$ and then $h:=k/d$ on the RHS of (1) yields
RHS $=\sum_{\begin{subarray}{c}i\\ i|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|i\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/i}=1\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{ij/e}}{2ij/e}$
$=\sum_{\begin{subarray}{c}d\\ d|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|(k/d)\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{k/(de)}=1\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{jd}}{2jd}$
$=\sum_{\begin{subarray}{c}h\\ h|k\end{subarray}}\sum_{\begin{subarray}{c}e\\ e|h\\ e|j\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=1\end{subarray}}K_{\varepsilon,\theta}(e)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}.$
@@
(2) $\sum_{\begin{subarray}{c}e\\ e|\gcd(h,j)\end{subarray}}\sum_{\begin{subarray}{c}\varepsilon\\ \varepsilon^{h/e}=1\end{subarray}}K_{\varepsilon,\theta}(e)=\begin{cases}\rho_{h}^{{\mathbb{Z}}_{2}}(\theta),&\text{if }\gcd(h,j)=1;\\ 0,&\text{otherwise.}\end{cases}$
@@
If $\gcd(h,j)\neq 1$ then the RHS here is zero, and if $\gcd(h,j)=1$ then the RHS is $\rho_{h}^{{\mathbb{Z}}_{2}}(\theta)$. This proves (2) and completes the proof of the theorem. ∎